OpenStudy (anonymous):

find the zeros of 2t^4+3t^3+16t-24

OpenStudy (anonymous):

\[ 16t-24=8(2t+3) \]And \[ 2t^4+3t^3=t^3(2t+3) \]

OpenStudy (anonymous):

So factoring gets you\[ (t^3+8)(2t+3) \]

OpenStudy (anonymous):

This is how you approach these problems. Finding the \(0\)s for the factors should not be hard.

OpenStudy (anonymous):

\[ 2t+3=0 \]And \[ t^3+8=0 \]

OpenStudy (anonymous):

but you have to make the \[t^3 +8 \] = something so would you just end up with 2 as the answer for that one?

OpenStudy (anonymous):

wait it would be -2

OpenStudy (anonymous):

Yeah